Find the real zeros of the polynomial function by factoring. The number in parentheses to the right of the polynomial indicates the number of real zeros of the given polynomial function. (Enter your answers as a comma-separated list.)
P(x) = x^3 − 3x^2 − 28x
(3)
x =

Answer by ankor@dixie-net.com(22740)

P(x) = x^3 − 3x^2 − 28x
Factor out x
x(x^2 - 3x - 28) = 0
factor the quadratic
x(x-7)(x+4) = 0
Three zeros
x = 0, 7, -4